OpenStudy (anonymous):

solve for x: log(3x-7)-log(x-2)=1

OpenStudy (amistre64):

might be a good idea to take the base of each side ...

OpenStudy (amistre64):

if log pertains to base 10, then 10 each side in higher math courses there is only 1 log, the natural log and therefore your notation would refer to a base e

OpenStudy (amistre64):

the idea is:\[b^{n+m}=b^n~b^m\]and \[b^1=b\] along with:\[b^{(k)_b}=k\]

OpenStudy (anonymous):

will it look like this 10^log(3x-7) - 10^log(x-2) = 1 3x-7-x-2=100 x=55 ?

OpenStudy (amistre64):

good try, but not quite: 10^ [log(3x-7) - log(x-2)] = 10^1 10^[log(3x-7)] * 10^[log(x-2)] = 10 (3x-7)(x-2) = 10

OpenStudy (amistre64):

even i messed that up lol, i dropped the negative

OpenStudy (amistre64):

10^ [log(3x-7) - log(x-2)] = 10^1 10^[log(3x-7)] * 10^[-log(x-2)] = 10 ^^ 10^[log(3x-7)] * 10^[log(1/(x-2))] = 10 (3x-7)/(x-2) = 10 thats better

OpenStudy (amistre64):

so the question becomes: 3x-2 = 10(x-2), solve for x

OpenStudy (amistre64):

gotta read thru the typos this morning .... 3x-7 = 10(x-2)

OpenStudy (anonymous):

why is it [-log(1/x-2)] ?

OpenStudy (amistre64):

by this property, and it went from -log(a) to log(1/a) \[k~log(a)=log(a^k)\] since k=-1, thats another way of presenting the reciprocal of a since:\[a^{-1}=1/a\]

OpenStudy (amistre64):

another way to approach it is this, same results tho: using the log property that:\[log(a/b)=log(a)-log(b)\] we can combine the logs, then base it

OpenStudy (anonymous):

that's exponential property or power property

OpenStudy (anonymous):

oh okay :)

OpenStudy (amistre64):

:) i cant recall what they name the things

OpenStudy (anonymous):

x= 13/7

OpenStudy (amistre64):

thats what i get

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=log%283x-7%29-log%28x-2%29%3D1&a=*FunClash.log-_*Log10- the wolf agrees, assuming log refers to base10